### v3

# Chaotic stochastic resonance in Mackey–Glass equations

Abstract
Stochastic resonance (SR) is observed as switching dynamics between two quasi-stationary states in stochastic Mackey–Glass equations. We identify a new form of SR, chaotic SR, characterized by positive Lyapunov exponents, arising from the coexistence of SR and stochastic chaos. Unlike stable SR, which exhibits negative Lyapunov exponents, the resonance point for chaotic SR precedes the zero-crossing point of the largest Lyapunov exponent. Furthermore, we provide a theoretical estimate of the resonant periods for both stable and chaotic SR based on a linear mode analysis around an unstable fixed point.
Stochastic resonance (SR) is a noise-induced phenomenon in which the response of a nonlinear system to an external force is enhanced by the addition of an optimal level of noise. SR can also arise in time-delayed systems, where the delayed feedback plays the role of an external force. In this study, we investigate the stochastic Mackey–Glass equation, a chaotic time-delayed system subject to noise, and identify two distinct types of SR, stable SR and chaotic SR. Since SR has often been examined from the perspective of stochastic processes, the underlying dynamical properties remain unclear. To clarify these properties, we apply random dynamical system theory to characterize the dynamical structure of the stochastic system. Stable SR is characterized by random point attractors, whereas chaotic SR is characterized by random strange attractors.
I Introduction
Adding noise to a deterministic dynamical system can induce qualitative changes in its behavior, known collectively as noise-induced phenomena. Stochastic resonance (SR) Gammaitoni et al. (1998) is a representative example, in which the system’s response to an external force is enhanced by an optimal level of noise. SR was originally investigated in bistable systems subjected to noise and periodic forcingBenzi, Sutera, and Vulpiani (1981); BENZI et al. (1982), where the addition of noise was found to enhance the periodicity of the system’s dynamics. Since then, SR has been observed in a broad range of nonlinear systems. Even without periodic forcing, noise-induced periodicity can emerge ( i.e., coherence resonance Pikovsky and Kurths (1997)). Likewise, the interplay between time delay and noise can generate a resonance in the absence of external forcing, known as delayed SR Ohira and Sato (1999); Tsimring and Pikovsky (2001); Masoller (2002). We refer to these phenomena collectively as SR in a broad sense. Based on the idea that deterministic chaos may act similarly to noise, related SR-like behavior in chaotic systems without external noise has also been studied, commonly referred to as chaotic resonance Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Krawiecki et al. (2001); Jüngling et al. (2008).
SR has also been reported in chaotic systems under both external noise and periodic forcing Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992). However, these studies have largely relied on the framework of stochastic processes, leaving the dynamical nature of resonance in chaotic regimes insufficiently clarified. To address this gap, we employ random dynamical system theory Arnold (1998); Chekroun, Simonnet, and Ghil (2011), which enables analysis of the intrinsic dynamical structure of stochastic systems. Within this framework, we identify stable SR characterized by random point attractors, and chaotic SR characterized by random strange attractors.
This paper is organized as follows. Section II.1 introduces the stochastic Mackey–Glass equation and its phenomenology. Section III presents numerical and analytical results for both stable and chaotic SR. Section IV provides concluding remarks.
II Model and phenomenology
II.1 Stochastic Mackey–Glass equations
Delay differential equations (DDEs) are widely used to model systems where time delay plays a crucial role Ikeda, Daido, and Akimoto (1980); Ikeda and Matsumoto (1987); Suarez and Schopf (1988); Keane, Krauskopf, and Postlethwaite (2017); Mackey and Glass (1977); Longtin and Milton (1988); Milton et al. (1989); Roxin, Brunel, and Hansel (2005). Despite their deceptively simple formulations, DDEs are infinite-dimensional dynamical systems capable of exhibiting rich behaviors, including high-dimensional chaos. The Mackey–Glass equation (MG) is a well-known DDE, developed as a model of hematopoiesis Mackey and Glass (1977), and is given by
$\frac{dx(t)}{dt}=\frac{ax(t-\tau)}{1+x^{c}(t-\tau)}-bx(t),$ (1)
where $\tau>0$ denotes the delay time, $a>0$ is the strength of delayed feedback, $b>0$ is the decay rate, and $c$ is the shape parameter of the delayed feedback. The initial condition is defined as $x(t)=\phi(t)$ for $t\in[-\tau,0)$ and is assumed to be constant unless otherwise specified. The MG can exhibit multiple positive Lyapunov exponents, and the effective dimension of the dynamics increases as $\tau$ grows Doyne Farmer (1982).
To investigate the impact of external noise, we introduce the stochastic Mackey–Glass equation (SMG) by incorporating additive Gaussian noise into Eq. (1):
$dx(t)=\Big[\frac{ax(t-\tau)}{1+x^{c}(t-\tau)}-bx(t)\Big]dt+\sigma dW_{t},$ (2)
where $\sigma$ represents the noise intensity and $W_{t}$ denotes a Wiener process. The parameters are fixed at the standard values $b=0.1$ and $c=10$, while $a$ and $\sigma$ serve as control parameters. The delay time is chosen to be sufficiently larger than the response time Mensour and Longtin (1998) $1/b$ and is set to $\tau=90$ unless stated otherwise.
II.2 Bifurcation in the stochastic Mackey–Glass equation
We briefly summarize the phenomenology of the dynamics in the MG and SMG. In the bifurcation diagram of the MG (Fig. 1 (a)), the Poincaré section, $dx/dt=0$, of the attractor in Eq. (1) is plotted as a function of the delayed feedback strength $a$ (blue dots). In the bifurcation diagram of the SMG (Fig. 1 (b)), projections of the random pullback attractors are shown as black dots (see Appendices A and B for details on pullback attractors and bifurcation diagram). The first and second Lyapunov exponents of the MG and SMG as functions of $a$ at $\sigma=0.0$ and $\sigma=0.15$ are shown in Fig. 1 (c) (blue and black, respectively). For the MG, the origin is stable for $0<a<0.1$. A pitchfork bifurcation occurs at $a=0.1$, giving rise to two symmetric fixed points with respect to the origin. As $a$ increases, a Hopf bifurcation emerges at $a\simeq 0.125$, producing a limit cycle. A bifurcation to chaos follows at $a\simeq 0.138$, with high-dimensional chaos appearing around $a\simeq 0.144$. In the SMG, the zero-crossing point of the first Lyapunov exponent occurs at $a=a_{c}\simeq 0.175$, beyond which stochastic chaosChekroun, Simonnet, and Ghil (2011) is observed. At $a=a_{d}\simeq 0.195$, the second Lyapunov exponent becomes positive, indicating the emergence of high-dimensional stochastic chaos.
In summary, although the MG becomes multi-stable via a pitchfork bifurcation, the SMG exhibits a single attractor with two quasi-stable states as two deterministic attractors are connected by external noise. As $a$ increases, the Lyapunov exponents of the SMG grow monotonically, leading to stochastic chaos and later to high-dimensional stochastic chaos. Notably, these stochastic bifurcations occur at significantly larger $a$ values than in the deterministic system.
Figure 1: (a) Bifurcation diagram of the MG with $\sigma=0$. (b) Bifurcation diagram of random pullback attractors of the SMG with $\sigma=0.15$. (c) First and second Lyapunov exponents for $\sigma=0$ (blue points) and $\sigma=0.15$ (black points). In the bifurcation diagram, random pullback attractors are computed using $1.0\times 10^{2}$ initial conditions with a pullback time $t_{p}=4\times 10^{4}$. We designate $a_{1}=0.14$ as a representative example of a weakly nonlinear regime and $a_{2}=0.30$ as that of a strongly nonlinear regime. At $a=a_{1},a_{2}$, low- and high-dimensional deterministic chaos are observed, respectively. The values, $a_{c}$ and $a_{d}$, indicate the zero-crossing points of the first and second Lyapunov exponents in the presence of noise. Numerical simulations are conducted using the Euler–Maruyama method with a time step of $\Delta t=0.1$, integrating for $10^{8}$ steps.
II.3 Stochastic resonance in the stochastic Mackey–Glass equation
Beyond the deterministic pitchfork bifurcation point ($a>0.1$), SR appears as switching dynamics between two quasi-stable states, associated with point, periodic, and strange attractors in the MG. Figure 2 illustrates SR for a regime with deterministic periodic attractors. Without noise, two symmetric limit cycles exist with respect to the origin (Fig. 2 (a)), each with a period of approximately $2\tau$. In the presence of noise, switching between these limit cycles emerges, and periodicity is enhanced at an optimal noise intensity. Figure 2 (c) shows the power spectra with and without noise (black and blue lines, respectively). Under optimal noise intensity, the peaks at $f=f_{n}^{*}\simeq n/\tau$ $(n=1,2,\ldots)$ reveal clear harmonics. The power of the primary resonant frequency $f_{1}^{*}$ reaches a maximum at the optimal noise intensity $\sigma=\sigma^{*}$, confirming the presence of SR. We refer to $\sigma^{*}$ as the resonance point (Fig. 2 (d)).
Figure 2: (a) Two deterministic attractors projected onto $(x(t),x(t-\tau))$ for $a=0.126$. The inset shows a magnified view of the limit cycle. (b) Time series for the deterministic case, $\sigma=0$ (blue dots), and the resonance case, $\sigma=\sigma^{*}$ (black dots). (c) Power spectra for $\sigma=0$ (blue lines) and the optimal noise intensity $\sigma=\sigma^{*}$ (black lines). (d) Power of the primary resonant frequency $f_{1}^{*}$ as a function of $\sigma$.
III Chaotic stochastic resonance
III.1 Stable and chaotic SR
To characterize the dynamical properties of SR, we compute the Lyapunov exponents of the SMG (see Appendix A for the definition of Lyapunov exponent in random dynamical systems). We examine two representative cases: one where the deterministic attractors exhibit low-dimensional chaos ($a=a_{1}=0.14$), and one where they exhibit high-dimensional chaos ($a=a_{2}=0.30$). In both cases, switching dynamics between two quasi-stable states appears, and the power spectrum peaks at $f^{*}_{n}\simeq n/\tau$ $(n=1,2,\ldots)$ are enhanced at the resonance point (Fig. 3). The sign of the largest Lyapunov exponent at the resonance point distinguishes the two types of SR: a negative exponent corresponds to stable SR (Fig. 3 (b)), whereas a positive exponent corresponds to chaotic SR (Fig. 3 (d)).
Figure 3: Power spectra for $\sigma=0$ (blue lines) and $\sigma=\sigma^{*}$ (black lines) at (a) $a=a_{1}$ and (c) $a=a_{2}$. Power of the primary resonant frequency $f_{1}^{*}$ (black points) as a function of $\sigma$ at (b) $a=a_{1}$ and (d) $a=a_{2}$. The largest Lyapunov exponent at $a=a_{1}$ and the first six Lyapunov exponents at $a=a_{2}$ are also shown in (b) and (d), respectively (red points).
In the case of stable SR, increasing the noise intensity causes $\lambda_{1}$ to become negative at $\sigma=\sigma_{0}$, at which point the chaotic attractor collapses into a random point attractor. As $\sigma$ is increased further, $\lambda_{1}$ attains a local maximum near the resonance point. Because the zero-crossing point $\sigma_{0}$ precedes the resonance point $\sigma^{*}$, chaos vanishes at resonance, rendering the system effectively analogous to the classical SR model with a double-well potential and periodic forcing Benzi, Sutera, and Vulpiani (1981). For sufficiently large $\sigma$, $\lambda_{1}$ decreases again, as trajectories stay almost always in contracting regions outside the attractors. In the case of chaotic SR, $\lambda_{1}$ initially increases and peaks near the resonance point. Here, $\sigma^{*}$ precedes $\sigma_{0}$, indicating that SR occurs while the system still possesses a positive Lyapunov exponent and thus retains stochastic chaos. Chaotic SR can be regarded as a class of noise-induced phenomena that enhances the characteristic periods of stochastic chaos. The ordering of $\sigma^{*}$ and $\sigma_{0}$ provide a clear diagnostic: $\sigma_{*}>\sigma_{0}$ for stable SR and $\sigma_{*}<\sigma_{0}$ for chaotic SR.
Stable SR, represented by $a=a_{1}$, is robustly observed in a weakly nonlinear regime $0.1<a<0.15$. Chaotic SR, represented by $a=a_{2}$, appears in a strongly nonlinear regime $a>0.28$, which includes deterministic window regions where noise-induced chaos Mayer-Kress and Haken (1981) arises. The occurrence of chaotic SR in such regimes suggests that high-dimensionality may contribute to its emergence, warranting further investigation.
III.2 Visualization of stable and chaotic SR
Previous studies on SR in chaotic systems have generally not distinguished chaotic SR from stable SR Anishchenko, Neiman, and Safanova (1993); Anishchenko, Safonova, and Chua (1992); Carroll and Pecora (1993); Nicolis, Nicolis, and McKernan (1993); Crisanti et al. (1994); Reibold et al. (1997); Krawiecki et al. (2001); Jüngling et al. (2008), largely due to the lack of a framework rooted in random dynamical system theory. The distinction becomes clear when visualizing the dynamics using space–time representations Arecchi et al. (1992); Yanchuk and Giacomelli (2017) and random pullback attractors (see Appendix A for the construction of random pullback attractors). A space–time representation maps the time series into a two-dimensional plane by introducing a discrete step $n\in\mathbb{Z}$ and a memory space $s\in[0,\tau)$. Both stable SR and chaotic SR appear as traveling waves in the memory space $s$ (Fig. 4, left column). Notably, chaotic SR exhibits pronounced periodicity, with chaotic structures clearly visible inside the traveling waves, indicating the coexistence of SR and stochastic chaos. When the largest Lyapunov exponent is negative, the random pullback attractor collapses to a random point attractor because initially close trajectories converge and eventually synchronize. Denoting a distribution of all possible trajectories at time $t$ as $\rho_{t}$, stable SR can be interpreted as pseudo-periodic motion of $\rho_{t}$ on the random point attractor (Fig.4 (b)). Conversely, when Lyapunov exponents are positive, the random pullback attractor becomes a random strange attractor. In this setting, chaotic SR corresponds to pseudo-periodic motion of $\rho_{t}$ on a random strange attractor (Fig. 4 (d)). This pseudo-periodic motion of $\rho_{t}$ is related to the problem of statistical periodicity Lasota and Mackey (1987); Losson and Mackey (1995). The precise relationship between chaotic SR and statistical periodicity remains an open question and presents an interesting direction for future study.
Figure 4: Space–time representation at the resonance point for (a) stable SR ($a=a_{1},\sigma=0.162$) and (c) chaotic SR ($a=a_{2},\sigma=0.104$). Both cases display traveling waves in memory space $s\in[0,\tau)$. The angle $\theta$ denotes the propagation direction of the traveling wave relative to the orthogonal direction, defined by $\tan\theta=\epsilon$. Using the measured angles $\theta_{1}$ (stable SR) and $\theta_{2}$ (chaotic SR), we obtain $\theta_{2}/\theta_{1}\simeq 0.786$, indicating that wave propagation is slower in the chaotic SR regime. (b) and (d) show snapshots of the random pullback attractor at the resonance point computed from $2.0\times 10^{4}$ initial conditions after a pullback time $t_{p}=2\times 10^{4}$, projected onto $(x(t),x(t-\tau))$.
III.3 Resonant period and unstable spiral
In studies of SR in DDEs, the resonant period $T=1/f^{*}_{1}$ is often approximated as $T\simeq\tau$. However, the traveling waves observed in the space–-time representation (Fig. 4, left column) indicate that the resonance period is slightly longer than $\tau$. We refined estimate, $T=\tau(1+\epsilon)$, based on the linear mode of the unstable spiral around the origin (Fig. 5, left column). This approximation improves upon the classical one in the regime where the delay $\tau$ is sufficiently larger than the response time $1/b$, but not excessively large. For sufficiently large $\tau$, the frequencies of the linear modes at $x=0$ are given by Mensour and Longtin (1998); Amann, Schöll, and Just (2007)
$s_{2m}=\frac{m}{\tau}\left[1-\frac{1}{b\tau+\ln(a\tau)}\right]~~(m=0,\pm 1,\cdots).$ (3)
We find that the $n$th resonant frequency satisfies $f^{*}_{n}\simeq s_{2n}~(n=1,2,\ldots)$, and thus the resonance period can be estimated as (see Appendix C)
$T=\frac{1}{f_{1}^{*}}\simeq\frac{1}{s_{2}}=\tau(1+\varepsilon),~~~~\varepsilon=\frac{1}{b\tau+\ln(a\tau)-1}.$ (4)
The classical approximation, $T\simeq\tau$, holds only when $\tau$ is extremely large. Numerical validation confirms that the primary resonant frequency, $f_{1}^{*}$ (black points), agrees well with the theoretical estimate, $s_{2}$ (red line), as shown in Figs. 5 (b) and (d).
Figure 5: Unstable spirals in the weakly and strongly nonlinear regimes. Phase-space portraits projected onto $(x(t),x(t-\tau))$ for (a) $a=a_{1}$ and (c) $a=a_{2}$. Transients initialized near the origin (red dots) eventually converge to other stable attractors (blue dots). The trajectory is computed from the initial function $\phi(t)=0.01\sin(2\pi f^{*}_{1}t)+0.02\sin(2\pi f^{*}_{2}t)$ and departs from the origin following a spiral structure. The primary resonant frequency $f_{1}^{*}$ as a function of $\tau$ for (b) $a=a_{1}$ and (d) $a=a_{2}$. The theoretical estimate (4) (red lines) agrees well with numerical results (black points). The classical estimate $f_{1}^{*}=1/\tau$ is also shown (dotted lines).
IV Conclusion
We investigated the dynamics of the Mackey–Glass equation in the presence of noise. In the weakly nonlinear regime, stable SR emerges as switching dynamics between two quasi-stationary states, with the largest Lyapunov exponent remaining negative at the resonance point. In this regime, the resonance point $\sigma^{*}$ follows the zero-crossing point of the largest Lyapunov exponent $\sigma_{0}$ ($\sigma_{0}<\sigma^{*}$). The deterministic chaotic attractor becomes stabilized into a random point attractor, and a resonance phenomenon in the classical double well potential model Benzi, Sutera, and Vulpiani (1981) effectively occurs. In the strongly nonlinear regime, we discovered chaotic SR, characterized by the coexistence of SR and stochastic chaos with positive Lyapunov exponents. Here, the resonance point $\sigma^{*}$ precedes the zero-crossing point $\sigma_{0}$ ($\sigma^{*}<\sigma_{0}$), indicating that the resonance occurs while the system remains chaotic. We also numerically observed the existence of chaotic SR in the stochastic Duffing equation. Under the overdamped limit, the stochastic Duffing equation reduces to the classical double well potential model, which exhibits stable SR. Thus, chaotic SR is expected to represent a universal class of noise-induced phenomena in strongly nonlinear random dynamical systems. Further investigation of noise-induced behavior in generic multi-attractor systems represents a promising avenue for future research.
Acknowledgements.
Authors thank Prof. T. Ohira (Nagoya University) for valuable comments. Y.S. is supported by JSPS Grant-in-Aid for Scientific Research (B), JP No. 21H01002.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Definitions of random pullback attractor and random Lyapunov exponent
To parametrize a noise realization, $\omega\in\Omega$, in time $t$, we introduce a family of measure-preserving maps, $\theta_{t}:\varOmega\rightarrow\varOmega$ satisfied with $\theta_{0}=\rm{id}_{\Omega}$ and $\theta_{s+t}=\theta_{s}\circ\theta_{t}$ for all $t$. The time evolution of random dynamical systems is described by a stochastic flow, $\varPhi(t,\omega):X\rightarrow X$, satisfying the cocycle property, $\varPhi(t+s,\omega)=\varPhi(t,\theta_{s}(\omega))\circ\varPhi(s,\omega)$. For systems with a single attractor, a set $\mathcal{A}(\omega)$ is defined as a random pullback attractor if it satisfies the following three conditions (see Arnold (1998); Chekroun, Simonnet, and Ghil (2011) for the exact definition):
  1. i.
    $\mathcal{A}(\omega)$ is a compact, i.e., $A(\omega):=\{x\in X|~(\omega,x)\in\mathcal{A}(\omega)\}\subset X$ is compact for almost all $\omega\in\Omega$.
  2. ii.
    $\mathcal{A(\omega)}$ is $\varPhi$-invariant, i.e., for all $t$, $\varPhi(t,\omega)A(\omega)=A(\theta_{t}\omega)$ for almost all $\omega\in\Omega$.
  3. iii.
    $\mathcal{A(\omega)}$ is pullback attracting, i.e., $\lim_{t\rightarrow\infty}d_{X}(\varPhi(t,\theta_{-t}\omega)B,A(\omega))=0$ holds for all $B\subset X$ and for almost all $\omega\in\Omega$.
where $d_{X}$ denotes the Hausdorff semi-distance.
A realization $A(\omega)$ of a random pullback attractor ${\mathcal{A}}(\omega)$ can be numerically approximated by a snapshot of trajectories evolving from a set of initial values $B$, with a fixed noise realization $\omega$, and with a integration period given by the pullback time $t_{p}$. For a precise numerical computation, a large number of initial points for $B$ and a large pullback time $t_{p}$ are adopted. The Lyapunov exponent of stochastic dynamics on random pullback attractors is given by the average expansion rate of perturbations, similarly to those of deterministic dynamics. For example, the Lyapunov exponent of a one-dimensional stochastic differential equation $dx=f(x)dt+\sigma dW_{t}$, where $W_{t}$ is the Wiener process, is given by
$\lambda(\omega,x_{0})=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}f^{\prime}(\varPhi(t,\omega)x_{0})dt,$ (5)
where $\varPhi$ represents the stochastic flow, and $x_{0}\in X$ is the initial condition. Note that the Lyapunov exponent is a random variable in general. When the system is ergodic and has a single attractor, $\lambda(\omega,x_{0})$ is a constant for almost all $\omega$ and for all $x_{0}$. As for high-dimensional random dynamical systems, the Lyapunov spectrum is defined in a similar manner to those of high-dimensional deterministic dynamical systems.
Appendix B Bifurcation diagram of random pullback attractors
We construct a bifurcation diagram of pullback attractors for the SMG. Plotting trajectories starting from many initial conditions with a fixed noise realization for each parameter, the bifurcation of random pullback attractors is successfully visualized. Similarly to random pullback attractors, this bifurcation diagram also evolves in time. In Fig. 6, emergence of the pseudo-periodic motion of the measure of all possible trajectories $\rho_{t}$ is observed near $a=0.14$ and $\sigma=0.15$.
Figure 6: Snapshots of bifurcation diagrams of random pullback attractors with a fixed noise realization for $\sigma=0.15$ and varying $a$. For all parameters, random pullback attractors are computed using a fixed noise realization. The numerical scheme follows that used in Fig. 1.
Appendix C Linear mode analysis of unstable spirals
When the delay $\tau$ is sufficiently large, the peaks of the power spectra of the Mackey–Glass equation (MG) agree with the linear modes around fixed points Mensour and Longtin (1998). The characteristic equations at fixed points $x=0,\pm x^{*}$, where $x^{*}=(\frac{a-b}{b})^{1/c}$
[…]
have an infinite number of roots $\chi=\lambda_{n},\mu_{n}~(n=0,\pm 1\pm 2,\ldots)$ given by
$\begin{array}[]{ll}\lambda_{n}=-b+\frac{1}{\tau}W_{n}\left[a\tau\exp(b\tau)\right]&(x=0),\\ \mu_{n}=-b+\frac{1}{\tau}W_{n}\left[(\frac{cb^{2}}{a}-(c-1)b)\tau\exp(b\tau)\right]&(x=\pm x^{*}),\end{array}$ (11)
where $W_{n}(z)$ is the Lambert W-function Corless et al. (1996). Using the asymptotic expansion of the Lambert W-function Amann, Schöll, and Just (2007), the frequencies of linear modes around the fixed points are given by the imaginary part of $\lambda_{n},\mu_{n}$, respectively expressed as
$\begin{array}[]{ll}s_{2n}&:=\frac{1}{2\pi}Im(\lambda_{n})\\ &\simeq\frac{2n}{2\tau}\left[1-\frac{1}{b\tau+\ln(a\tau)}\right]~~~~~~~~~~~~~~~~~~~~~~~~(x=0),\\ s_{2n+1}&:=\frac{1}{2\pi}Im(\mu_{n})\\ &\simeq\frac{2n+1}{2\tau}\left[1-\frac{1}{b\tau+\ln(|\frac{cb^{2}}{a}-(c-1)b|\tau)}\right]~~~~~(x=\pm x^{*}).\end{array}$ (12)
The even modes $s_{2n}$ represent an unstable spiral around the origin, and the odd modes $s_{2n+1}$ represent spirals around the fixed point $x=\pm x^{*}$. The characteristic frequencies of the dynamics dominated by odd modes at $x=\pm x^{*}$ (Fig. 7). Adding the optimal noise, the odd modes are suppressed and the even modes become dominant, causing the emergence of stochastic resonance (SR) with resonant frequencies corresponding to the even mode.
Figure 7: (a) Roots of linear modes $\chi=\lambda_{n}$ (black) and $chi=\mu_{n}$ (blue) at $a=a_{1}=0.14$. (b) Power spectra for $a=a_{1}=0.14$ with $\sigma=0$ (blue line) and optimal noise $\sigma=\sigma^{*}$ (black line).